# Log Parsing with Self-Generated In-Context Learning and Self-Correction

### v1

@@ Abstract.
Log parsing transforms log messages into structured formats, serving as a crucial step for log analysis. Despite a variety of log parsing methods that have been proposed, their performance on evolving log data remains unsatisfactory due to reliance on human-crafted rules or learning-based models with limited training data. The recent emergence of large language models (LLMs) has demonstrated strong abilities in understanding natural language and code, making it promising to apply LLMs for log parsing. Consequently, several studies have proposed LLM-based log parsers. However, LLMs may produce inaccurate templates, and existing LLM-based log parsers directly use the template generated by the LLM as the parsing result, hindering the accuracy of log parsing. Furthermore, these log parsers depend heavily on historical log data as demonstrations, which poses challenges in maintaining accuracy when dealing with scarce historical log data or evolving log data. To address these challenges, we propose AdaParser, an effective and adaptive log parsing framework using LLMs with self-generated in-context learning (SG-ICL) and self-correction. To facilitate accurate log parsing, AdaParser incorporates a novel component, a template corrector, which utilizes the LLM to correct potential parsing errors in the templates it generates. In addition, AdaParser maintains a dynamic candidate set composed of previously generated templates as demonstrations to adapt evolving log data. Extensive experiments on public large-scale datasets show that AdaParser outperforms state-of-the-art methods across all metrics, even in zero-shot scenarios. Moreover, when integrated with different LLMs, AdaParser consistently enhances the performance of the utilized LLMs by a large margin.
1. Introduction
@@
Log messages are generated by logging statements in the source code to record system execution behaviors. By inspecting system logs, engineers can detect anomalies (Du et al., 2017), localize software bugs (Chen et al., 2021), or troubleshoot problems (Jia et al., 2021) in the system. Therefore, log analysis plays a crucial role in software maintenance. The prerequisite step of log analysis is log parsing (He et al., 2021), which extracts two parts of log messages: (1) log templates. i.e., constant parts that are explicitly written in logging statements, and (2) variables, dynamic parts that vary during runtime and reflect system runtime status.
@@
Manual log parsing is impractical due to the sheer volume of logs. Hence, numerous automatic log parsers have been proposed (Zhang et al., 2023), which can be categorized into syntax-based and semantic-based log parsers. Syntax-based log parsers (Fu et al., 2009; Nagappan and Vouk, 2010; He et al., 2017; Dai et al., 2020) use human-crafted features or heuristics (e.g., log length and frequency) to extract log templates. In contrast, semantic-based log parsers (Liu et al., 2022; Le and Zhang, 2023b; Huo et al., 2023b; Li et al., 2023) employ deep learning models to mine semantics from log data for parsing. However, recent studies have revealed that the performance of existing log parsers deteriorates over time when handling evolving log data (Wang et al., 2022; Yu et al., 2023c; Liu et al., 2023b). On the one hand, syntax-based log parsers, which rely heavily on predefined rules, experience significant performance degradation when processing logs that deviate from these rules. On the other hand, semantic-based log parsers, which typically employ deep learning models, struggle to adapt to evolving log data without periodic retraining, which is constrained by the quantity and quality of labeled data available.
@@
To address these limitations, we intend to leverage the large language models (LLMs) to achieve adaptive log parsing for evolving log data. LLMs are trained on extensive corpora and exhibit strong capabilities in understanding natural language and code (Zhao et al., 2023; Fan et al., 2023), making them well-suited for understanding log messages. Furthermore, LLM’s powerful in-context learning (ICL) ability allows them to perform downstream tasks without model tuning (Dong et al., 2022; Liu et al., 2023c). Although several studies (Le and Zhang, 2023a; JIANG et al., 2023; Liu et al., 2023b; Xu et al., 2024b; Ma et al., 2024) have explored using LLMs for log parsing, designing an effective and adaptive LLM-based log parser for evolving log data still faces the following challenges:
@@
First, despite their strong text understanding and generation capabilities, LLMs may produce inaccurate parsing results. For example, the template generated by the LLM fails to accurately match the input log message or is overly broad, i.e., identifying crucial information as variables, which is essential for troubleshooting. Second, existing LLM-based parsers (JIANG et al., 2023; Xu et al., 2024b) heavily rely on historical log data to sample candidates as demonstrations. However, frequent software updates often result in evolving log data, making sufficient historical log data unavailable. Our experiments indicate that these parsers perform well with abundant historical log data but struggle to maintain accuracy when historical log data is limited.
@@
We have conducted extensive experiments to evaluate AdaParser on public large-scale log datasets (Jiang et al., 2023). The results show that AdaParser achieves a 96.5% F1 score in grouping accuracy and an 87.9% F1 score in template accuracy, outperforming state-of-the-art baselines and exhibiting remarkable robustness across diverse log datasets. Furthermore, AdaParser surpasses all existing baselines in accuracy across various ratios of available logs, even in the absence of historical logs, underscoring its adaptability. When integrated with different LLMs, AdaParser consistently enhances its performance on all metrics by a large margin. In terms of efficiency, AdaParser achieves a speed comparable to the most efficient baseline, Drain (He et al., 2017). The evaluation results demonstrate that AdaParser is an effective, adaptive, and efficient log parsing framework suitable for deployment in real-world evolving systems.
In summary, this paper makes the following main contributions:
@@
  • We propose an effective and adaptive LLM-based log parsing framework named AdaParser. With the proposed self-generated ICL and self-correction, AdaParser can perform accurate log parsing and adapt to evolving log data.
@@
  • We evaluate AdaParser on public large-scale datasets. The results show that AdaParser outperforms state-of-the-art methods across all metrics, even in zero-shot scenarios. Additionally, when integrated with different LLMs, AdaParser consistently enhances its performance by a large margin.
Figure 1. An example of log parsing from Spark.
2. background and motivation
2.1. Log Parsing
@@
Log parsing is a fundamental step in log analysis (He et al., 2021; Zhang et al., 2023), converting semi-structured log messages to structured data. As illustrated in Figure 1, in log parsing, a log parser is used to extract log headers (e.g., datetime “17/08/22 15:51:02”, verbosity level “INFO”, and component name “BlockManager”) and log contents (e.g., “BlockManager Writing block rdd_1_3 to disk”). Since log headers are automatically generated by the logger, their format is fixed and can usually be extracted easily. Therefore, log parsers primarily focus on extracting the constant parts (i.e., log templates) and the dynamic parts (i.e., variables) from log contents. A straightforward method is to match log messages with corresponding logging statements within the source code (Pecchia et al., 2015; Schipper et al., 2019). However, this method is impractical when the source code is inaccessible, such as commercial software. Consequently, a variety of data-driven log parsers without requiring access to the source code have been proposed (He et al., 2017; Le and Zhang, 2023b; Yu et al., 2023a), which can be categorized into syntax-based and semantic-based methods. Syntax-based log parsers (Fu et al., 2009; Nagappan and Vouk, 2010; He et al., 2017; Dai et al., 2020) use human-crafted features or heuristics (e.g., log length and frequency) to extract log templates. In contrast, semantic-based log parsers (Liu et al., 2022; Le and Zhang, 2023b; Huo et al., 2023b; Li et al., 2023) employ deep learning models to mine semantics from log data for parsing.
@@
In modern software systems, software updates lead to continuous changes in logging statements, resulting in evolving log data (Zhang et al., 2019; Liu et al., 2023a; Huo et al., 2023a). For instance, Google has reported thousands of newly added logging statements each month due to software updates (Xu, 2010). Recent studies have indicated that existing log parsers’ performance degrades over time when handling evolving log data (Wang et al., 2022; Yu et al., 2023c; Liu et al., 2023b). On the one hand, syntax-based log parsers heavily rely on human-crafted rules, requiring a substantial amount of domain-specific knowledge. This limitation results in poor performance when processing evolving log data that do not conform to these rules. For example, Drain (He et al., 2017), a leading syntax-based log parser, assumes that logs from the same template have the same length, which can lead to parsing errors when logs belonging to the same template vary in length. On the other hand, semantic-based log parsers, which typically adopt deep learning models, struggle to adapt to evolving log data without periodic retraining, which is inherently limited by the quantity and quality of labeled data available for model tuning and involves expensive tuning processes. Therefore, there is a need for a log parser that does not rely on human-crafted rules and can adapt to evolving log data without requiring model tuning.
2.2. Large Language Models
@@
Large Language Models (LLMs) are large-sized pre-trained language models with tens or hundreds of billions of parameters, trained on extensive unlabeled corpora via self-supervised learning. LLMs exhibit strong capabilities in understanding natural language (Zhao et al., 2023). Beyond natural language, LLMs are also effective in understanding code, which has spurred interest in their application within the software engineering domain (Fan et al., 2023). Effectively applying LLMs to downstream tasks is an important research topic. A common approach involves fine-tuning the model by learning the desired output from the given input of a downstream dataset to update model parameters (Wang et al., 2021). However, fine-tuning can be extremely time-consuming and resource-intensive, especially for models with billions of parameters. Moreover, the performance of fine-tuned models heavily relies on the scale and quality of labeled data, making it less feasible in scenarios with limited labeled data.
@@
Recently, in-context learning (ICL) has emerged as a promising alternative that enables LLMs to perform downstream tasks without the necessity for explicit model tuning (Dong et al., 2022; Liu et al., 2023c). In the ICL paradigm, a formatted natural language prompt is used to query LLMs. This prompt typically includes three parts: (1) Instruction, which describes the specific task; (2) Demonstrations, which consist of several examples (i.e., query-answer pairs) selected from the labeled task datasets; and (3) Query, which is the actual question that LLMs need to answer. The number of demonstrations included in the prompt can vary, defining different scenarios. Specifically, if there are no demonstrations, the scenario is known as zero-shot learning; if there is only one demonstration, the scenario is known as one-shot learning; and few-shot learning means there are several demonstrations. Numerous studies have demonstrated that ICL can significantly enhance the performance of LLMs across diverse tasks, such as log statement generation (Xu et al., 2024a), comment generation (Geng et al., 2024), and code review (Guo et al., 2024). Given these advancements, we intend to apply the ICL paradigm to enable adaptive log parsing for evolving log data, capitalizing on the ability of LLMs to learn and adapt from structured examples without requiring continuous retraining.
2.3. Challenges of Existing LLM-based parsers
@@
While recent advancements (Le and Zhang, 2023a; JIANG et al., 2023; Liu et al., 2023b; Xu et al., 2024b; Ma et al., 2024) have shown the potential of using LLMs for log parsing, these LLM-based parsers face the following critical challenges that limit their broader application.
2.3.1. Unsatisfactory accuracy in parsing logs
@@
Despite their strong text understanding and generating capabilities, LLMs often produce inaccurate parsing results. Figure 2 illustrates two major types of parsing errors from LILAC (JIANG et al., 2023), a leading LLM-based parser: (1) Plausible template, which occurs when the template generated by the LLM fails to accurately match the input log message (e.g., regular expression matching). Specifically, the LLM incorrectly parsed “0x300sent” as a variable “0x300” and a constant “sent”. such errors reduce grouping accuracy and negatively impact downstream tasks like log compression and anomaly detection. (2) Broad template, which happens when the generated template matches the log message but is overly broad. Specifically, the LLM identified the exception message “Could not read from stream” as a variable. However, this exception message is crucial for troubleshooting and should be preserved in the log template. Such broad templates degrade template accuracy, affecting downstream tasks such as root cause analysis. Therefore, mitigating these parsing errors in the template generated by the LLM remains an unresolved challenge.
2.3.2. Poor adaptability to evolving logs
@@
Existing LLM-based parsers (Xu et al., 2024b; JIANG et al., 2023) rely heavily on historical log data to sample candidates as demonstrations. However, in practice, software updates frequently result in evolving logs, making sufficient historical log data unavailable. In extreme cases, such as when a brand-new service goes online, no historical log data exists. Our experiments also indicate that these parsers perform well with sufficient historical log data but struggle to maintain accuracy when such data is limited. For example, the template accuracy of LILAC decreases significantly by 31.2% when in the absence of historical log data. Consequently, enabling LLM-based log parsers to adapt to evolving logs without relying on extensive historical data is a substantial challenge.
@@
In summary, while LLM-based log parsers present significant potential, several challenges must be addressed to enhance their accuracy and adaptability, which is essential to fully harness the potential of LLMs in log parsing and develop solutions that are both practical and effective in real-world applications.
Figure 2. Two types of parsing errors in LLM-based parser.
Figure 3. The overview of AdaParser framework.
3. AdaParser
3.1. Overview
@@
For each log message, AdaParser first employs the tree-based parser to check whether its corresponding template is already stored in the parsing tree. If a match is found, AdaParser directly uses the matched template as the parsed result of this log message, which avoids redundant LLM queries and improves parsing efficiency. Otherwise, Adaparser utilizes the LLM to generate a new template for this log message. To mitigate the degradation of parsing accuracy due to evolving log data, we propose an adaptive SG-ICL-based Parser. This parser maintains a dynamic candidate set using previously LLM-generated templates and selects demonstrations from the candidate set to construct the parsing prompt, guiding the LLM to parse logs. To reduce potential parsing errors in LLM-generated templates, we propose a novel and effective template corrector to correct the template. The corrected template is then used to update both the candidate set and the parsing tree.
3.2. Tree-based Log Parser

### v2

# Log Parsing using LLMs with Self-Generated In-Context Learning and Self-Correction

@@ Abstract
Log parsing transforms log messages into structured formats, serving as a crucial step for log analysis. Despite a variety of log parsers that have been proposed, their performance on evolving log data remains unsatisfactory due to reliance on human-crafted rules or learning-based models with limited training data. The recent emergence of large language models (LLMs) has demonstrated strong abilities in understanding natural language and code, making it promising to apply LLMs for log parsing. Consequently, several studies have proposed LLM-based log parsers. However, LLMs may produce inaccurate templates, and existing LLM-based log parsers directly use the template generated by the LLM as the parsing result, hindering the accuracy of log parsing. Furthermore, these log parsers depend heavily on historical log data as demonstrations, which poses challenges in maintaining accuracy when dealing with scarce historical log data or evolving log data. To address these challenges, we propose AdaParser, an effective and adaptive log parsing framework using LLMs with self-generated in-context learning (SG-ICL) and self-correction. To facilitate accurate log parsing, AdaParser incorporates a novel component, a template corrector, which utilizes the LLM to correct potential parsing errors in the templates it generates. In addition, AdaParser maintains a dynamic candidate set composed of previously generated templates as demonstrations to adapt evolving log data. Extensive experiments on public large-scale datasets indicate that AdaParser outperforms state-of-the-art methods across all metrics, even in zero-shot scenarios. Moreover, when integrated with different LLMs, AdaParser consistently enhances the performance of the utilized LLMs by a large margin.
Index Terms:
Log Parsing, Large Language Model, In-Context Learning, Self-Correction
I Introduction
@@
Log messages are generated by logging statements in the source code to record system runtime information [1, 2, 3]. By inspecting system logs, engineers can detect anomalies [4], localize software bugs [5], or troubleshoot problems [6] in the system. Therefore, log analysis plays a crucial role in software maintenance. The first step of log analysis is log parsing [7], which extracts two parts of log messages: (1) log templates. i.e., constant parts that are explicitly written in logging statements, and (2) variables, i.e., dynamic parts that vary during runtime and reflect system runtime status.
@@
Manual log parsing is impractical due to the sheer volume of logs. Hence, numerous automatic log parsers have been proposed [8], which can be categorized into syntax-based and semantic-based log parsers. Syntax-based log parsers [9, 10, 11, 12] use human-crafted features or heuristics (e.g., log length and frequency) to extract log templates. In contrast, semantic-based log parsers [13, 14, 15, 16] employ deep learning models to mine semantics from log data for parsing. However, recent studies have revealed that the performance of existing log parsers declines over time when handling evolving log data [17, 18, 19]. On the one hand, syntax-based log parsers, which rely heavily on predefined rules, experience significant performance degradation when processing logs that deviate from these rules. On the other hand, semantic-based log parsers, which typically employ deep learning models, struggle to adapt to evolving log data without periodic retraining, which is constrained by the quantity and quality of labeled data available.
@@
To address these limitations, we intend to leverage the large language models (LLMs) to achieve adaptive log parsing for evolving log data. LLMs are trained on extensive corpora and exhibit strong capabilities in understanding natural language and code [20, 21], making them well-suited for understanding log messages. Furthermore, LLM’s powerful in-context learning (ICL) ability allows them to perform downstream tasks without model tuning [22, 23]. Although several studies [19, 24, 25, 26, 27] have explored using LLMs for log parsing, designing an effective and adaptive LLM-based log parser for evolving log data still faces the following challenges:
@@
First, despite their strong text understanding and generation abilities, LLMs may produce inaccurate parsing results, and existing LLM-based log parsers directly use the template generated by the LLM as the parsing result, hindering the accuracy of log parsing. For example, the template generated by the LLM fails to accurately match the input log message or is overly broad, i.e., identifying crucial information as variables, which is essential for troubleshooting.
Second, existing LLM-based parsers [25, 26] heavily rely on historical log data to sample candidates as demonstrations. However, software updates often result in evolving log data, making sufficient historical log data unavailable. Our experiments indicate that these parsers perform well with abundant historical log data but struggle to maintain accuracy when historical log data is limited.
@@
We have conducted extensive experiments to evaluate AdaParser on public large-scale log datasets [28]. The results show that AdaParser achieves a 96.5% F1 score in grouping accuracy and an 87.9% F1 score in template accuracy, outperforming state-of-the-art baselines and exhibiting remarkable robustness across diverse log datasets. Furthermore, AdaParser surpasses all existing baselines in accuracy across various ratios of available logs, even in the absence of historical logs, underscoring its adaptability. When integrated with different LLMs, AdaParser consistently enhances its performance on all metrics by a large margin. In terms of efficiency, AdaParser achieves a speed comparable to the most efficient baseline, Drain [11]. The evaluation results demonstrate that AdaParser is an effective, adaptive, and efficient log parsing framework suitable for deployment in real-world evolving systems.
In summary, this paper makes the following contributions:
@@
  • We propose AdaParser, an effective and adaptive LLM-based log parsing framework. With the proposed self-generated ICL and self-correction, AdaParser can perform accurate log parsing and adapt to evolving log data.
@@
  • We evaluate AdaParser on public large-scale datasets. The results show that AdaParser outperforms state-of-the-art methods on all metrics, even in zero-shot scenarios. Additionally, when integrated with different LLMs, AdaParser consistently enhances its performance by a large margin.
  • We release the code of AdaParser at https://github.com/wuyifan18/AdaParser to benefit both practitioners and researchers in the field of log analysis.
Figure 1: An example of log parsing from Spark.
II background and motivation
II-A Log Parsing
@@
Log parsing is a fundamental step in log analysis [7, 8], converting semi-structured log messages to structured data. As illustrated in Figure 1, in log parsing, a log parser is used to extract log headers (e.g., datetime “17/08/22 15:51:02”, verbosity level “INFO”, and component name “BlockManager”) and log contents (e.g., “BlockManager Writing block rdd_1_3 to disk”). Since log headers are automatically generated by the logger, their format is fixed and can usually be extracted easily. Therefore, log parsers primarily focus on extracting the constant parts (i.e., log templates) and the dynamic parts (i.e., variables) from log contents. A straightforward method is to match log messages with corresponding logging statements within the source code [29, 30]. However, this method is impractical when the source code is inaccessible, such as commercial software. Consequently, a variety of data-driven log parsers without requiring access to the source code have been proposed [11, 14, 31], which can be categorized into syntax-based and semantic-based methods. Syntax-based log parsers [9, 10, 11, 12] use human-crafted features or heuristics (e.g., log length and frequency) to extract log templates. In contrast, semantic-based log parsers [13, 14, 15, 16] employ deep learning models to mine semantics from log data for parsing.
@@
In modern software systems, software updates lead to continuous changes in logging statements, resulting in evolving log data [32, 33, 34]. For instance, Google has reported thousands of newly added logging statements each month due to software updates [35]. Recent studies have indicated that existing log parsers’ performance degrades over time when handling evolving log data [17, 18, 19]. On the one hand, syntax-based log parsers heavily rely on human-crafted rules, requiring a substantial amount of domain-specific knowledge. This limitation results in poor performance when processing evolving log data that do not conform to these rules. For example, Drain [11], a leading syntax-based log parser, assumes that logs from the same template have the same length, which can lead to parsing errors when logs belonging to the same template vary in length. On the other hand, semantic-based log parsers, which typically adopt deep learning models, struggle to adapt to evolving log data without periodic retraining, which is inherently limited by the quantity and quality of labeled data available for model tuning and involves expensive tuning processes. Therefore, there is a need for a log parser that does not rely on human-crafted rules and can adapt to evolving log data without requiring model tuning.
II-B Large Language Models
@@
Large Language Models (LLMs) are large-sized pre-trained language models with tens or hundreds of billions of parameters, trained on extensive unlabeled corpora via self-supervised learning. LLMs exhibit strong capabilities in understanding natural language [20]. Beyond natural language, LLMs are also effective in understanding code, which has spurred interest in their application within the software engineering domain [21]. Effectively applying LLMs to downstream tasks is an important research topic. A common approach involves fine-tuning the model by learning the desired output from the given input of a downstream dataset to update model parameters [36]. However, fine-tuning can be extremely time-consuming and resource-intensive, especially for models with billions of parameters. Moreover, the performance of fine-tuned models heavily relies on the scale and quality of labeled data, making it less feasible in scenarios with limited labeled data.
@@
Recently, in-context learning (ICL) has emerged as a promising alternative that enables LLMs to perform downstream tasks without the necessity for explicit model tuning [22, 23]. In the ICL paradigm, a formatted natural language prompt is used to query LLMs. This prompt typically includes three parts: (1) Instruction, which describes the specific task; (2) Demonstrations, which consist of several examples (i.e., query-answer pairs) selected from the labeled task datasets; and (3) Query, which is the actual question that LLMs need to answer. The number of demonstrations included in the prompt can vary, defining different scenarios. Specifically, if there are no demonstrations, the scenario is known as zero-shot learning; if there is only one demonstration, the scenario is known as one-shot learning; and few-shot learning means there are several demonstrations. Numerous studies have demonstrated that ICL can significantly enhance the performance of LLMs across diverse tasks, such as log statement generation [1], comment generation [37], and code review [38]. Given these advancements, we intend to apply the ICL paradigm to enable adaptive log parsing for evolving log data, capitalizing on the ability of LLMs to learn and adapt from structured examples without requiring continuous retraining.
II-C Challenges of Existing LLM-based parsers
@@
While recent advancements [24, 25, 19, 26, 27] have shown the potential of using LLMs for log parsing, these LLM-based parsers face the following critical challenges that limit their broader application.
II-C1 Unsatisfactory accuracy in parsing logs
@@
Despite their strong text understanding and generating capabilities, LLMs often produce inaccurate parsing results. Figure 2 illustrates two major types of parsing errors from LILAC [25], a leading LLM-based parser: (1) Plausible template, which occurs when the template generated by the LLM fails to accurately match the input log message (e.g., regular expression matching). Specifically, the LLM incorrectly parsed “0x300sent” as a variable “0x300” and a constant “sent”. such errors reduce grouping accuracy and negatively impact downstream tasks like log compression and anomaly detection. (2) Broad template, which happens when the generated template matches the log message but is overly broad. Specifically, the LLM identified the exception message “Could not read from stream” as a variable. However, this exception message is crucial for troubleshooting and should be preserved in the log template. Such broad templates degrade template accuracy, affecting downstream tasks such as root cause analysis. Therefore, mitigating these parsing errors in the template generated by the LLM remains an unresolved challenge.
II-C2 Poor adaptability to evolving logs
@@
Existing LLM-based parsers [26, 25] rely heavily on historical log data to sample candidates as demonstrations. However, in practice, frequent software updates result in evolving logs, making sufficient historical log data unavailable. In extreme cases, such as when a brand-new service goes online, no historical log data exists. Our experiments also indicate that these parsers perform well with sufficient historical log data but struggle to maintain accuracy when such data is limited. For example, the template accuracy of LILAC decreases significantly by 31.2% when in the absence of historical log data. Consequently, enabling LLM-based log parsers to adapt to evolving logs without relying on extensive historical data is a substantial challenge.
@@
In summary, while LLM-based log parsers show significant potential, several challenges must be addressed to enhance their accuracy and adaptability, which is essential to fully harness the potential of LLMs in log parsing and develop practical and effective solutions for real-world applications.
Figure 2: Two types of parsing errors in LLM-based parser.
Figure 3: The overview of AdaParser framework.
III Approach
III-A Overview
@@
For each log message, AdaParser first employs the tree-based parser to check whether its corresponding template is already stored in the parsing tree. If a match is found, AdaParser directly uses the matched template as the parsed result of this log message, which avoids redundant LLM queries and improves parsing efficiency. Otherwise, Adaparser utilizes the LLM to generate a new template for this log message. To mitigate the degradation of parsing accuracy due to evolving log data, we utilize the SG-ICL-based Parser to maintain a dynamic candidate set using previously LLM-generated templates and select demonstrations from the candidate set to construct the parsing prompt, guiding the LLM to parse logs. To reduce potential parsing errors in LLM-generated templates, we propose a novel and effective template corrector to correct the template. The corrected template is then used to update both the candidate set and the parsing tree.
III-B Tree-based Log Parser